# Notions of enriched purity

### v2

@@ 4. A weaker notion
Let $f\colon K\to L$ be a $(\lambda,\mathcal{E})$-pure morphism in $\mathcal{K}$ and $g\colon A\to B$ a morphism in $\mathcal{K}_{\lambda}$. The universal property of the pullback defining $\mathcal{P}(g,f)$ induces a map $t\colon\mathcal{K}^{\to}(g,f)\to\mathcal{P}(g,f)$ making the square below commute
@@
and, since $p^{\prime}\in\mathcal{E}$ there exists $s$ in $\mathcal{M}$ as above making the triangles commute. By pre-composition with $q^{\prime}$, we obtain that $sq=r$; thus, if $(\mathcal{E},\mathcal{M})$ is proper, then $s$ is also in $\mathcal{E}$ and hence is an isomorphism. Therefore $q$ is also in $\mathcal{E}$ and $f$ is barely $(\lambda,\mathcal{E})$-pure. Assume now that $\mathcal{E}$ is stable under pullbacks in $\mathcal{V}$. The map $t$ can be seen as the pullback of the morphism $\mathcal{K}(B,L)\twoheadrightarrow\mathcal{P}(g,L)$ along $\mathcal{P}(g,f)\to\mathcal{P}(g,L)$; thus $t$ is in $\mathcal{E}$ by hypothesis. The uniqueness of the $(\mathcal{E},\mathcal{M})$-factorization then implies that $s\colon\mathcal{Q}(g,f)\to\mathcal{P}(g,f)$ above is an isomorphism. Thus $f$ is $(\lambda,\mathcal{E})$-pure is and only if it is barely $(\lambda,\mathcal{E})$-pure. ∎
@@
where $g$ is from 3.4(3) and is dense. But $\emptyset\to\{0\}$ is not dense. Nevertheless, $(\lambda,\mathcal{E})$-pure and barely $(\lambda,\mathcal{E})$-pure are equivalent notions by 6.10.
@@
If $\mathcal{E}\subseteq\mathcal{E}^{\prime}$ then $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ and thus there exists $t:\mathcal{Q}_{\mathcal{E}}(g,f)\to\mathcal{Q}_{\mathcal{E}^{\prime}}(g,f)$ such that $tp^{\prime}_{\mathcal{E}}=p^{\prime}_{\mathcal{E}^{\prime}}$. Following [9, 2.1.4], $t\in\mathcal{E}^{\prime}$ and thus
$q_{\mathcal{E}^{\prime}}=p^{\prime}_{\mathcal{E}^{\prime}}q^{\prime}=tq_{\mathcal{E}}q^{\prime}=tq_{\mathcal{E}}$
belongs to $\mathcal{E}^{\prime}$. ∎
@@
If $(\mathcal{E},\mathcal{M})$ is proper then barely $(\lambda,\mathcal{E})$-pure morphisms are left-cancellable; i.e., if $f_{2}f_{1}$ is barely $(\lambda,\mathcal{E})$-pure then $f_{1}$ is barely $(\lambda,\mathcal{E})$-pure.
@@
Let $f_{2}f_{1}$ be $(\lambda,\mathcal{E})$-pure. Let
@@
be the $(\mathcal{E},\mathcal{M})$ factorizations where $f=f_{2}f_{1}$. Since $(\operatorname{id}_{K},f_{2}):f_{1}\to f$, we have
$\mathcal{K}^{\to}(g,(\operatorname{id}_{K},f_{2})):\mathcal{K}^{\to}(g,f_{1})\to\mathcal{K}^{\to}(g,f)$
and $p\mathcal{K}^{\to}(g,(\operatorname{id}_{K},f_{2}))=p_{1}$. Thus there is $s:\mathcal{Q}(g,f_{1})\to\mathcal{Q}(g,f)$ such that
@@
and $p^{\prime\prime}s=p^{\prime\prime}_{1}$. Moreover, $sq_{1}=q$ and thus $s$ is in $\mathcal{E}$. Since it is in $\mathcal{M}$, $s$ is an isomorphism. Hence $q_{1}$ is in $\mathcal{E}$ and thus $f_{1}$ is $(\lambda,\mathcal{E})$-pure. ∎
@@
Same arguments of Corollary 3.11, now relying on the proposition above. ∎
@@
Since $\mathcal{M}$ is closed under limits in $\mathcal{V}^{\to}$, it is enough to prove the result for $C\in\mathcal{K}_{\lambda}$. Furthermore, by Corollary 4.11 above, it is enough to show that any barely $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ in $\mathcal{V}$ lies in $\mathcal{M}$. By the hypotheses on $\mathcal{M}$, to prove that $f\in\mathcal{M}$ it is enough to show that it is right orthogonal to every morphism in $\mathcal{E}_{\lambda}$; for that it suffices to prove that the square below is a pullback.
@@
The notion of $\mathcal{E}$-split monomorphism should not be confused with that of $\mathcal{E}$-split epimorphism; this is a map $s$ for which $\mathcal{K}(L,s)\colon\mathcal{K}(L,K)\to\mathcal{K}(K,K)$ is required to be in $\mathcal{E}$.
@@
    If $\mathcal{E}$ is the class of surjections in the sense of 3.3 then $\mathcal{E}$-split monomorphisms are precisely the usual split monomorphisms.
@@
Remark 4.19.
The same results on change of base, given in Remark 3.20, hold with the same hypothesis if we replace $(\lambda,\mathcal{E})$-purity with bare $(\lambda,\mathcal{E})$-purity.
@@ 5. The characterization theorem
    where $\omega+1$ is the successor of $\omega$ seen as an ordinal (this is obtained from $\omega$ by freely adding the join of all elements), $y$ sends the only element of $1$ to the top element of $\omega+1$, and $d$ sends the countable antichain $\omega\cdot 1$ to the chain $\omega\subseteq\omega+1$.
@@
$(i)\Rightarrow(ii)$. By definition of bare $(\lambda,\mathcal{E})$-purity and since $\mathcal{K}$ is locally $\lambda$-presentable, it is enough to choose $\mathcal{D}$ to be any $\lambda$-filtered family of $\lambda$-presentable objects of $\mathcal{K}^{\to}$ with colimit $f$. $(ii)\Rightarrow(i)$. Fix a family $\mathcal{D}=\{f_{i}\colon K_{i}\to L_{i}\}_{i\in I}$ as in condition $(ii)$ and consider any $g\colon A\to B$ between $\lambda$-presentable objects. Then $\mathcal{K}(B,K)\cong\operatorname{colim}_{i}\mathcal{K}(B,K_{i})$, $\mathcal{K}^{\to}(g,f)\cong\operatorname{colim}_{i}\mathcal{K}^{\to}(g,f_{i})$, and (since $\mathcal{M}$ and $\mathcal{E}$ are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$) also $\mathcal{Q}(g,f)\cong\operatorname{colim}_{i}\mathcal{Q}(g,f_{i})$ and $q\cong\operatorname{colim}q_{i}$ (with the notation of 4.1). We need to show that $q\colon\mathcal{K}(B,K)\to\mathcal{Q}(g,f)$ lies in $\mathcal{E}$. For that, write $\mathcal{E}$ as an injectivity class $\mathcal{H}$-inj in $\mathcal{V}^{\to}$ as in the initial assumption. Thus, we are required to show that $q$ is injective with respect to each map $(!,y)\colon!_{X}\to d$ in $\mathcal{H}$. Consider then a morphism $(!,v)\colon!_{X}\to q$, since $X$ is $\lambda$-presentable and $\mathcal{Q}(g,f)$ is the $\lambda$-filtered colimit of the $\mathcal{Q}(g,f_{i})$ it follows that $v$ factors as a map $v_{i}\colon X\to\mathcal{Q}(g,f_{i})$, for a fixed $i$. We can then construct the following commutative diagram:
@@
Since $\mathcal{E}$ is an injectivity class, it is closed in $\mathcal{V}^{\to}$ under products. Thus the necessity follows from 3.17.
Conversely, let $\mathcal{A}$ be closed under products, $\lambda$-filtered colimits, and $(\lambda,\mathcal{E})$-pure subobjects in $\mathcal{K}$. Since $\mathcal{A}$ is closed also under $\lambda$-pure subobjects by Corollary 3.9, $\mathcal{A}_{0}$ is a $\lambda$-injectivity class of $\mathcal{K}_{0}$ by [27, Theorem 2.2]. Thanks to [16, Theorem 3.3] we then obtain a functor $R\colon\mathcal{K}_{0}\to\mathcal{K}_{0}$ preserving $\lambda$-filtered colimits and a natural transformation $r\colon 1_{\mathcal{K}_{0}}\to R$ such that each component $r_{K}\colon K\to RK$ is an ordinary weak reflection of $K$ into $\mathcal{A}_{0}$.
@@
The fact that $\mathcal{A}\subseteq\textnormal{Inj}\mathcal{H}$ is true by definition. Consider then $K\in\textnormal{Inj}\mathcal{H}$ together with its ordinary weak reflection $r_{k}\colon K\to RK$ into $\mathcal{A}$. To conclude it is enough to show that $r_{K}$ is $(\lambda,\mathcal{E})$-pure, so that $K\in\mathcal{A}$. For this, we use Lemma 5.3.
Write $K\cong\operatorname{colim}K_{i}$ as a $\lambda$-filtered colimit of $\lambda$-presentable objects in $\mathcal{K}$ with colimiting maps $x_{i}\colon K_{i}\to K$; since $R$ preserves $\lambda$-filtered colimits then $RK\cong\operatorname{colim}RK_{i}$ and $r_{K}\cong\operatorname{colim}r_{K_{i}}$ in $\mathcal{K}^{\to}$. Now, by writing each $RK_{i}$ as a $\lambda$-filtered colimit $s_{j}:L_{j}\to RK_{i}$ of $\lambda$-presentable objects, we get that each $r_{K_{i}}$ is a $\lambda$-filtered colimit of morphisms $h_{j}:K_{i}\to L_{j}$ in $K_{i}\downarrow\mathcal{K}$ via $s_{j}:L_{j}\to R_{K_{i}}$. It follows that the family $\mathcal{D}:=\{h_{j}:K_{i}\to L_{j}\}_{i,j}$ is $\lambda$-directed, contained in $\mathcal{K}_{\lambda}^{\to}$, and has colimit $r_{K}$. By Lemma 5.3, to conclude it is enough to show that $r_{K}$ is $\mathcal{E}$-pure with respect to each $h_{j}$.
Since $r_{K_{i}}=s_{j}h_{j}$ and $r_{K_{i}}$ is an ordinary weak reflection, every object from $\mathcal{A}$ is injective in the ordinary sense with respect to $h_{j}$; that is $\mathcal{A}_{0}(A,h_{j})$ is surjective for any $A\in\mathcal{A}$. Since by hypothesis $\mathcal{A}$ is closed under powers by $\mathcal{G}$, it follows that for any $A\in\mathcal{A}$ and $G\in\mathcal{G}$ the maps
$\mathcal{V}_{0}(G,\mathcal{A}(A,h_{j}))\cong\mathcal{A}_{0}(G\pitchfork A,h_{j})$
are surjective. By $(ii)$ this means that $\mathcal{A}(A,h_{j})\in\mathcal{E}$, and thus every object from $\mathcal{A}$ is $\mathcal{E}$-injective with respect to $h_{j}$. Therefore $K$ is $\mathcal{E}$-injective with respect to $h_{j}$, i.e., $\mathcal{K}(h_{j},K)$ is in $\mathcal{E}$. Now, following Notation 4.1 for $g:=h_{j}$ and $f=r_{K}$ we obtain
$\mathcal{K}(h_{j},K)=pq^{\prime}=p^{\prime\prime}p^{\prime}q^{\prime}=p^{\prime\prime}q,$
thus $p^{\prime\prime}$ is in $\mathcal{E}$. Hence $p^{\prime\prime}$ is an isomorphism, therefore $q$ is in $\mathcal{E}$. Consequently, $r_{K}$ is $\mathcal{E}$-pure with respect to $h_{j}$, as claimed. ∎
@@
Let $\mathcal{V}=\operatorname{\bf Ban}$ with $(\mathcal{E},\mathcal{M})=$ (dense, isometry); then: $\mathcal{E}$ satisfies Assumptions 5.1, the two notions of purity coincide (by Example 4.20), and $\mathcal{G}:=\{\mathbb{C}\}$ satisfies condition (ii) above. Therefore we can apply Theorem 5.5 and obtain the following characterization: the $(\lambda,\mathcal{E})$-injectivity classes in a locally $\lambda$-presentable $\operatorname{\bf Ban}$-category $\mathcal{K}$ are precisely the classes closed under $\lambda$-filtered colimits, products, and $(\lambda,\mathcal{E})$-pure subobjects.
@@
Points (1) and (2) follows from Propositions 2.6 and 2.7 of [15]. For (3), assume that $f\colon K\to L$ is $(\lambda,\mathcal{E})$-pure and that $L\in\mathcal{H}^{\perp}$. For any $g\colon A\to B$ in $\mathcal{H}$, following Notation 3.1, we obtain that $\mathcal{P}(g,L)\cong\mathcal{K}(A,L)$ (since $\mathcal{K}(g,L)$ is an isomorphism), and hence that $\mathcal{P}(g,f)\cong\mathcal{K}(A,K)$. Thus, the $\mathcal{E}$-purity of $f$ with respect to $g$ implies that $\mathcal{K}(g,K)$ lies in $\mathcal{E}$. But by Proposition 3.12 we know that $\mathcal{K}(B,f)$ and $\mathcal{K}(A,f)$ are in $\mathcal{M}$, so that $\mathcal{K}(g,K)$ is in $\mathcal{M}$ too. It follows that $\mathcal{K}(g,K)$ is an isomorphism and hence that $K\in\mathcal{H}^{\perp}$. ∎

### v3

@@ 4. A weaker notion
Let $f\colon K\to L$ be a $(\lambda,\mathcal{E})$-pure morphism in $\mathcal{K}$ and $g\colon A\to B$ a morphism in $\mathcal{K}_{\lambda}$. In the notation of 4.1, the universal property of the pullback defining $\mathcal{P}(g,f)$ induces a map $t\colon\mathcal{K}^{\to}(g,f)\to\mathcal{P}(g,f)$ making the square below commute
@@
and, since $p^{\prime}\in\mathcal{E}$ there exists $s$ in $\mathcal{M}$ as dashed above making the triangles commute. By pre-composition with $q^{\prime}$, we obtain that $sq=r$, and $r$ is in $\mathcal{E}$ by $(\lambda,\mathcal{E})$-purity of $f$. Thus, if $(\mathcal{E},\mathcal{M})$ is proper, then $s$ is also in $\mathcal{E}$ and hence is an isomorphism. Therefore $q$ is also in $\mathcal{E}$ and $f$ is barely $(\lambda,\mathcal{E})$-pure.
Assume now that $\mathcal{E}$ is stable under pullbacks in $\mathcal{V}$. The map $t$ can be seen as the pullback of the morphism $\mathcal{K}(B,L)\twoheadrightarrow\mathcal{P}(g,L)$ along $\mathcal{P}(g,f)\to\mathcal{P}(g,L)$ from Notation 3.1; thus $t$ is in $\mathcal{E}$ by hypothesis. The uniqueness of the $(\mathcal{E},\mathcal{M})$-factorization then implies that $s\colon\mathcal{Q}(g,f)\to\mathcal{P}(g,f)$ above is an isomorphism. Thus $f$ is $(\lambda,\mathcal{E})$-pure if and only if it is barely $(\lambda,\mathcal{E})$-pure. ∎
@@
where $g$ is from 3.5(3) and is dense. But $\emptyset\to\{0\}$ is not dense.
Nevertheless, $(\lambda,\mathcal{E})$-pure and barely $(\lambda,\mathcal{E})$-pure are equivalent notions by 6.9.
@@
We will follow Notation 4.1 with indices denoting the factorization system to which they apply. If $\mathcal{E}\subseteq\mathcal{E}^{\prime}$ then $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ and thus there exists $t:\mathcal{Q}_{\mathcal{E}}(g,f)\to\mathcal{Q}_{\mathcal{E}^{\prime}}(g,f)$ such that $tp^{\prime}_{\mathcal{E}}=p^{\prime}_{\mathcal{E}^{\prime}}$. By properness of the factorization system, $t\in\mathcal{E}^{\prime}$ and thus
$q_{\mathcal{E}^{\prime}}=p^{\prime}_{\mathcal{E}^{\prime}}q^{\prime}=tp^{\prime}_{\mathcal{E}}q^{\prime}=tq_{\mathcal{E}}$
belongs to $\mathcal{E}^{\prime}$ (note that $q^{\prime}$ does not depend on $\mathcal{E}$ or $\mathcal{E}^{\prime}$). ∎
@@
If $(\mathcal{E},\mathcal{M})$ is proper then barely $(\lambda,\mathcal{E})$-pure morphisms are left-cancellable; that is, if $f_{2}f_{1}$ is barely $(\lambda,\mathcal{E})$-pure then $f_{1}$ is barely $(\lambda,\mathcal{E})$-pure.
@@
We follow Notation 4.1. Let $f_{2}f_{1}$ be $(\lambda,\mathcal{E})$-pure and consider the factorizations
@@
where $f=f_{2}f_{1}$. Since $(\operatorname{id}_{K},f_{2}):f_{1}\to f$ defines a morphism in $\mathcal{K}^{\to}$, we have a map
$\mathcal{K}^{\to}(g,(\operatorname{id}_{K},f_{2}))\colon\mathcal{K}^{\to}(g,f_{1})\to\mathcal{K}^{\to}(g,f)$
for which $p\mathcal{K}^{\to}(g,(\operatorname{id}_{K},f_{2}))=p_{1}$. Thus there is an induced $s\colon\mathcal{Q}(g,f_{1})\to\mathcal{Q}(g,f)$ such that
@@
and $p^{\prime\prime}s=p^{\prime\prime}_{1}$. Moreover, $sq_{1}=q$ and thus $s$ is in $\mathcal{E}$. Since it is also in $\mathcal{M}$, $s$ is an isomorphism. Hence $q_{1}$ is in $\mathcal{E}$ and thus $f_{1}$ is $(\lambda,\mathcal{E})$-pure. ∎
@@
Same arguments of Corollary 3.12, now relying on Proposition 4.10 above. ∎
@@
Since $\mathcal{M}$ is closed under limits in $\mathcal{V}^{\to}$, it is enough to prove the result for $C\in\mathcal{K}_{\lambda}$. Furthermore, by Corollary 4.11 above, it is enough to show that any barely $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ in $\mathcal{V}$ lies in $\mathcal{M}$.
By the hypotheses on $\mathcal{M}$, to prove that $f\in\mathcal{M}$ it is enough to show that it is right orthogonal to every morphism in $\mathcal{E}_{\lambda}$; for that it suffices to prove that the square below is a pullback.
@@
The notion of $\mathcal{E}$-split monomorphism should not be confused with that of $\mathcal{E}$-split epimorphism; this is a map $s$ for which $\mathcal{K}(L,s)\colon\mathcal{K}(L,K)\to\mathcal{K}(L,L)$ is required to be in $\mathcal{E}$.
@@
    If $\mathcal{E}$ is the class of surjections in the sense of 3.4 then $\mathcal{E}$-split monomorphisms are precisely the usual split monomorphisms.
@@ 5. The characterization theorem
    where $\omega+1$ is the successor of $\omega$ seen as an ordinal (this is obtained from $\omega$ by freely adding the join of all elements), $y$ sends the only element of $1$ to the top element of $\omega+1$, and $d$ sends the countable anti-chain $\omega\cdot 1$ to the chain $\omega\subseteq\omega+1$.
@@
$(i)\Rightarrow(ii)$. By definition of bare $(\lambda,\mathcal{E})$-purity and since $\mathcal{K}$ is locally $\lambda$-presentable, it is enough to choose $\mathcal{D}$ to be any $\lambda$-filtered family of $\lambda$-presentable objects of $\mathcal{K}^{\to}$ with colimit $f$.
$(ii)\Rightarrow(i)$. Fix a family $\mathcal{D}=\{f_{i}\colon K_{i}\to L_{i}\}_{i\in I}$ as in condition $(ii)$ and consider any $g\colon A\to B$ between $\lambda$-presentable objects. We will use notation 4.1 with indices $i$’s denoting the $f_{i}$ to which they apply. Then $\mathcal{K}(B,K)\cong\operatorname{colim}_{i}\mathcal{K}(B,K_{i})$ and $\mathcal{K}^{\to}(g,f)\cong\operatorname{colim}_{i}\mathcal{K}^{\to}(g,f_{i})$. Moreover, since $\mathcal{M}$ and $\mathcal{E}$ are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$, also $\mathcal{Q}(g,f)\cong\operatorname{colim}_{i}\mathcal{Q}(g,f_{i})$ and $q\cong\operatorname{colim}q_{i}$. We need to show that $q\colon\mathcal{K}(B,K)\to\mathcal{Q}(g,f)$ lies in $\mathcal{E}$.
For that, write $\mathcal{E}$ as an injectivity class $\mathcal{H}$-inj in $\mathcal{V}^{\to}$ as in the initial assumption. Thus, we are required to show that $q$ is injective with respect to each map $(!,y)\colon!_{X}\to d$ in $\mathcal{H}$.
Consider then a morphism $(!,v)\colon!_{X}\to q$, since $X$ is $\lambda$-presentable and $\mathcal{Q}(g,f)$ is the $\lambda$-filtered colimit of the $\mathcal{Q}(g,f_{i})$ it follows that $v$ factors as a map $v_{i}\colon X\to\mathcal{Q}(g,f_{i})$, for a fixed $i$. We can then construct the following commutative diagram:
@@
where $M$ is induced by first applying the projection $\mathcal{K}^{\to}(g,f_{i})\to\mathcal{K}(B,L_{i})$ and then composing, and similarly $N$ is induced through the factorization system by the composition maps in $\mathcal{K}$ and $\mathcal{K}^{\to}$. Finally, the map $v^{\prime}$ is given by tensoring the composite of the colimiting map $I\to\mathcal{K}(f_{i},f)$ and $p_{i}^{\prime}\colon\mathcal{K}(f_{i},f)\to\mathcal{Q}(f_{i},f)$ with $v_{i}$ (and composing with the isomorphism $X\cong I\otimes X$).
@@
Since $\mathcal{E}$ is an injectivity class, it is closed in $\mathcal{V}^{\to}$ under products. Thus the necessity follows from 3.18.
Conversely, let $\mathcal{A}$ be closed under products, $\lambda$-filtered colimits, and $(\lambda,\mathcal{E})$-pure subobjects in $\mathcal{K}$. Since $\mathcal{A}$ is closed also under $\lambda$-pure subobjects by Corollary 3.10, its underlying category $\mathcal{A}_{0}$ is a $\lambda$-injectivity class of $\mathcal{K}_{0}$ by [27, Theorem 2.2]. Thanks to [16, Theorem 3.3] we then obtain a functor $R\colon\mathcal{K}_{0}\to\mathcal{K}_{0}$ preserving $\lambda$-filtered colimits and a natural transformation $r\colon 1_{\mathcal{K}_{0}}\to R$ such that each component $r_{K}\colon K\to RK$ is an ordinary weak reflection of $K$ into $\mathcal{A}_{0}$.
@@
The fact that $\mathcal{A}\subseteq\textnormal{Inj}\mathcal{H}$ is true by definition. Consider then $K\in\textnormal{Inj}\mathcal{H}$ together with its ordinary weak reflection $r_{k}\colon K\to RK$ into $\mathcal{A}$. To conclude it is enough to show that $r_{K}$ is $(\lambda,\mathcal{E})$-pure, so that $K\in\mathcal{A}$ by the assumptions on $\mathcal{A}$. For this, we use Lemma 5.3.
Write $K\cong\operatorname{colim}K_{i}$ as a $\lambda$-filtered colimit of $\lambda$-presentable objects in $\mathcal{K}$ with colimiting maps $x_{i}\colon K_{i}\to K$; since $R$ preserves $\lambda$-filtered colimits then $RK\cong\operatorname{colim}RK_{i}$ and $r_{K}\cong\operatorname{colim}r_{K_{i}}$ in the category of arrows $\mathcal{K}^{\to}$. Now, by writing each $RK_{i}$ as a $\lambda$-filtered colimit
$s_{j}:L^{i}_{j}\to RK_{i}$
of $\lambda$-presentable objects, we get that each $r_{K_{i}}$ is a $\lambda$-filtered colimit of morphisms $h^{i}_{j}:K_{i}\to L^{i}_{j}$ in $K_{i}\downarrow\mathcal{K}$ via $s^{i}_{j}:L^{i}_{j}\to RK_{i}$. It follows that the family
$\mathcal{D}:=\{h^{i}_{j}:K_{i}\to L^{i}_{j}\}_{i,j}$
is $\lambda$-directed, contained in $\mathcal{K}_{\lambda}^{\to}$, and has colimit $r_{K}$. By Lemma 5.3, to conclude it is enough to show that $r_{K}$ is $\mathcal{E}$-pure with respect to each $h_{j}$.
Since $r_{K_{i}}=s_{j}h^{i}_{j}$ and $r_{K_{i}}$ is an ordinary weak reflection, every object from $\mathcal{A}$ is injective in the ordinary sense with respect to $h^{i}_{j}$; that is $\mathcal{A}_{0}(A,h^{i}_{j})$ is surjective for any $A\in\mathcal{A}$. Since by hypothesis $\mathcal{A}$ is closed under powers by $\mathcal{G}$, it follows that for any $A\in\mathcal{A}$ and $G\in\mathcal{G}$ the maps
$\mathcal{V}_{0}(G,\mathcal{A}(A,h^{i}_{j}))\cong\mathcal{A}_{0}(G\pitchfork A,h^{i}_{j})$
[…]
$\mathcal{K}(B,K)$$\mathcal{K}(A,K)$$\mathcal{Q}(g,f)$$\mathcal{K}(h^{i}_{j},K)$$q$$\ \ \ p^{\prime\prime}$
where $p^{\prime\prime}$ is in $\mathcal{M}$. Since $\mathcal{K}(h^{i}_{j},K)$ is in $\mathcal{E}$, also $p^{\prime\prime}$ is in $\mathcal{E}$. Hence $p^{\prime\prime}$ is an isomorphism, and therefore $q$ is in $\mathcal{E}$. Consequently, $r_{K}$ is $\mathcal{E}$-pure with respect to $h^{i}_{j}$, as claimed. ∎
@@
Let $\mathcal{V}=\operatorname{\bf Ban}$ with $(\mathcal{E},\mathcal{M})=$ (dense, isometry); then: $\mathcal{E}$ satisfies Assumptions 5.1, the two notions of purity coincide (by Example 4.19), and $\mathcal{G}:=\{\mathbb{C}\}$ satisfies condition (ii) above.
Therefore we can apply Theorem 5.5 and obtain the following characterization: the $(\lambda,\mathcal{E})$-injectivity classes in a locally $\lambda$-presentable $\operatorname{\bf Ban}$-category $\mathcal{K}$ are precisely the classes closed under $\lambda$-filtered colimits, products, and $(\lambda,\mathcal{E})$-pure subobjects.
@@
Points (1) and (2) follows from Propositions 2.6 and 2.7 of [15].
For (3), assume that $f\colon K\to L$ is $(\lambda,\mathcal{E})$-pure and that $L\in\mathcal{H}^{\perp}$. For any $g\colon A\to B$ in $\mathcal{H}$, following Notation 3.1, we obtain that $\mathcal{P}(g,L)\cong\mathcal{K}(A,L)$ (since $\mathcal{K}(g,L)$ is an isomorphism), and hence that $\mathcal{P}(g,f)\cong\mathcal{K}(A,K)$. Thus, the $\mathcal{E}$-purity of $f$ with respect to $g$ implies that $\mathcal{K}(g,K)$ lies in $\mathcal{E}$. But by Proposition 3.13 we know that $\mathcal{K}(B,f)$ and $\mathcal{K}(A,f)$ are in $\mathcal{M}$, so that $\mathcal{K}(g,K)$ is in $\mathcal{M}$ too. It follows that $\mathcal{K}(g,K)$ is an isomorphism and hence that $K\in\mathcal{H}^{\perp}$. ∎